A body of mass 0.2kg is executing shmwith an amplitude of 2pcm. The maximum force which acts upon it is 0.064N. Calculate it's maximum velocity and period of oscillation.

1
Expert's answer
2022-03-03T12:05:38-0500

Answer

Force

"F=mAw^2=0.064N"

So

"Aw^2=0.32"

maximum velocity

"V=Aw=\sqrt{Aw^2 A}\\=\sqrt{0.32*0.02}\\=0.08m/s"

Period of oscillation


"T=\frac{2\pi}{w}=\pi /2"